Find the sample mean x for each possible sample of size n = 2. Sample 4, 6 4, 8 4, 10 SampleX 6, 8 6, 10 8, 10 T. (Type integers or decimals rounded to two decimal places as needed.) 2
Answer #1

\mathbf{Solution}

\mathbf{Mean\;(\bar x)}=\frac{\sum x}{n}

\mathbf{Sample } \mathbf{\bar x} \mathbf{Sample } \mathbf{\bar x}
4,6 \frac{4+6}{2}={\color{Red} 5} 6,8 \frac{6+8}{2}={\color{Red} 7}
4,8 \frac{4+8}{2}={\color{Red} 6} 6,10 \frac{6+10}{2}={\color{Red} 8}
4,10 \frac{4+10}{2}={\color{Red} 7} 8,10 \frac{8+10}{2}={\color{Red} 9}